# SoundMorpher: Perceptually-Uniform Sound Morphing with Diffusion Model

### v1

@@ 4 Method
Given a source and target audio pair $\{x^{(0)},x^{(1)}\}$, sound morphing aims to generate intermediate sounds $x^{(\alpha(t))}$ between the audio pair given morph factors $\alpha\in[0,1]$. To account for the variation of $\alpha$ over time in Equation 1, we discretize the function $\alpha(t)$ where $t\in[0,T]$ into $N$ elements, resulting in a morphed sequence of sounds $\{x^{(\alpha_{i})}\}_{i=1}^{N}$ based on $\{\alpha_{i}\}^{N}_{i=1}$. According to the smoothness criteria proposed by Caetano & Osaka (2012), the desired sound morphing technique should have smooth linear perceptual stimuli when the morph factor $\alpha$ varies in the sequence $\{\alpha_{i}\}^{N}_{i=1}$. Therefore, we define $p_{i}$ to represent the perceptual stimuli of the morphed audio $x^{(\alpha_{i})}$ given morph factor $\alpha_{i}$. However, the relationship $\mathcal{P(\cdot)}$ between morph factor $\alpha$ and perceptual stimuli $p$ is intractable. Our goal is to find a discrete morph factor sequence $\{\alpha_{i}\}^{N}_{i=1}$ such that for each transition, the perceptual stimuli difference $\Delta p$ is a constant value. Therefore, we formulate the problem as
$p_{i+1}-p_{i}\equiv\mathcal{P}(x^{(\alpha_{i+1})})-\mathcal{P}(x^{(\alpha_{i})})=\Delta p,~{}~{}i\in[1,...,N-1]~{}$ (5)
@@
This formulation is a refined sound morphing problem where, rather than controlling morph factor $\alpha$, we control the constant perceptual stimuli difference $\Delta p$ to find the optimal trajectory with morph factors $\{\alpha_{i}\}^{N}_{i=1}$ that will achieve perceptually uniform sound morphing  [^1] See Appendix 7.1 for overall SoundMorpher pseudo algorithm pipeline and further implementation details..
@@
In Section 4.1 we introduce feature interpolation and model adaption with a pre-trained AudioLDM2. This method allows high-quality intermediate morph results to be obtained by controlling morph factor $\alpha$. To achieve perceptually uniform sound morphing as in Equation 5, we explore an explicit connection $\mathcal{P(\cdot)}$ between perceptual stimuli $p$ and morph factor $\alpha$ in Section 4.2. In Section 4.3, we provide extensions of our method on the different morphing methods discussed in Section 3.1 to show the advantages of perceptually uniform sound morphing.
4.1 Feature interpolation and model adaption
@@
Interpolating optimized conditional embeddings. We first introduce text-guided conditional embedding optimization strategy under a pre-trained AudioLDM2, which retrieves corresponding conditional embeddings $E$ of the given audio data. As mentioned in Section 3.2, AudioLDM2 accepts two conditional inputs: LOA feature $Y$ and text embedding $E_{T5}$. We denote $E=\{Y,E_{T5}\}$ as the overall conditional embedding inputs for AudioLDM2. The LOA feature $Y$ is an abstraction of audio data which is semantically structured, and $E_{T5}$ captures sentence-level of representations. To retrieve corresponding conditional embeddings of the given audio data, we first obtain the latent variables $z_{0}^{(0)}$ and $z_{0}^{(1)}$ of audio $x^{(0)}$ and $x^{(1)}$ from the pre-trained VAE in AudioLDM2 pipeline. We initialize a simple common text prompt (e.g., ‘An audio clip of sound’) as a text guidance condition $C$ to obtain $E$ by GPT-2 encoder and FLAN-T5 encoder in AudioLDM2 pipeline, respectively, as $E^{(0)}$ and $E^{(1)}$. Instead of optimizing the model parameters, we freeze the model parameters and optimize the conditional embedding $E^{(0)}$ and $E^{(1)}$ by the denoising objective function in Equation 2
$E^{(0)}=\text{arg min}_{E}\mathcal{L}_{DPM}(z_{0}^{(0)},E;\theta)\text{ and }E^{(1)}=\text{arg min}_{E}\mathcal{L}_{DPM}(z_{0}^{(1)},E;\theta)$ (6)
@@
Interpolating latent state. The conditional embedding represents the conceptual abstract of audio data. However, we also wish to smoothly morph the content of the audio pair. Following Song et al. (2020) and Yang et al. (2023), we smoothly interpolate between $z_{0}^{(0)}$ and $z_{0}^{(1)}$ by spherical linear interpolation (slerp) to their starting noise $z_{T}^{(0)}$ and $z_{T}^{(1)}$ and further obtained the interpolated latent state $z_{T}^{(\alpha)}:=\frac{\text{sin}(1-\alpha)\omega}{\text{sin}\omega}z_{T}^{(0)}+\frac{\text{sin}\alpha\omega}{\text{sin}\omega}z_{T}^{(1)}$, where $\omega=\text{arccos}(\frac{z_{T}^{(0)\intercal}z_{T}^{(1)}}{||z_{T}^{(0)}||||z_{T}^{(1)}||})$. The denoised latent variable $z_{0}^{(\alpha)}$ is obtained by applying a diffusion denoising process on the interpolated starting noise $z_{T}^{(\alpha)}$ and conditioning on the interpolated conditional embedding $E^{(\alpha)}$. The final morphed audio result $x^{(\alpha)}$ is obtained from $z_{0}^{(\alpha)}$ by the VAE decoder and a vocoder.
@@
Model adaption. Model adaptation helps to limit the degree of morphed variation by suppressing high-density regions that not related to the given inputs (Yang et al., 2023). We use LoRA (Hu et al., 2021) to inject a small amount of trainable parameters for efficient model adaptation. We fine-tune AudioLDM2 with LoRA trainable parameters using $z^{(0)}$ and $z^{(1)}$. See Appendix 7.2 for details.
4.2 perceptually uniform Sound Morphing
@@
Sound perceptual distance proportion (SPDP). The relationship between morph factor $\alpha$ and perceptual stimuli $p$ is intractable. Our goal is to establish an objective quantitative metric that links $p_{i}$ and $x^{(\alpha_{i})}$ as in Equation 5. This metric should satisfy two key conditions: (1) the output $p$ should increase monotonically as $\alpha$ increases; (2) it should accurately represent perceptual differences between $x^{(\alpha)}$ and $\{x^{(0)},x^{(1)}\}$, ensuring a smooth transition through intermediate states. Therefore, we propose the sound perceptual distance proportion between $x^{(\alpha)}$ and $\{x^{(0)},x^{(1)}\}$. We define $p_{i}\in\mathbb{R}^{2}$ as a 2D vector to represent the perceptual proximity of $x^{(\alpha_{i})}$ to both $x^{(0)}$ and $x^{(1)}$. Instead of extracting numerous audio features through traditional signal processing techniques, we use Mel-scaled spectrogram to capture perceptual and semantic information on audio. Mel-spectrogram (Tzanetakis & Cook, 2002) provides a pseudo-3D representation of audio signals, with one axis representing time and the other representing frequency on the Mel scale (Stevens et al., 1937), while the values denote the magnitude of each frequency at specific time points. The advantage of using Mel-spectrogram lies in the Mel filter banks, which map frequencies to equal pitch distances that correspond to how humans perceive sound (Sturm, 2013; Müller, 2015). Denoting $x_{mel}^{(\alpha_{i})}$ as the Mel-spectrogram of audio $x^{(\alpha_{i})}$, the SPDP $p_{i}$ between two endpoint audios $x^{(0)}$ and $x^{(1)}$ given $\alpha_{i}$ is defined as
$p_{i}=[\frac{||x_{mel}^{(\alpha_{i})}-x_{mel}^{(0)}||_{2}}{||x_{mel}^{(\alpha_{i})}-x_{mel}^{(0)}||_{2}+||x_{mel}^{(\alpha_{i})}-x_{mel}^{(1)}||_{2}},\frac{||x_{mel}^{(\alpha_{i})}-x_{mel}^{(1)}||_{2}}{||x_{mel}^{(\alpha_{i})}-x_{mel}^{(0)}||_{2}+||x_{mel}^{(\alpha_{i})}-x_{mel}^{(1)}||_{2}}]$ (7)
@@
Binary search with constant SPDP increment. To produce a perceptually smooth morphing trajectory with a constant perceptual stimuli increment, we use binary search to seek the corresponding $\{\alpha_{i}\}^{N}_{i=1}$ based on a constant $\Delta p$. The target SPDP sequence $\{p_{i}\}_{i=1}^{N}$ is obtained by an interpolation $p_{i}=(1-\frac{i-1}{N-1})p^{(0)}+\frac{i-1}{N-1}p^{(1)}$, where the two endpoints are $p^{(0)}=[0,1]^{T}$ and $p^{(1)}=[1,0]^{T}$. See Algorithm 2 in Appendix 7.3 for detail pseudo algorithm.
4.3 Controllable Sound Morphing with Discrete $\alpha$ Series
By controlling the discrete morph factor sequence $\{\alpha_{i}\}^{N}_{i=1}$ to produce a morphed series $\{x^{(\alpha_{i})}\}_{i=1}^{N}$ with constant $\Delta p$, we achieve three typical morphing methods as follows.

### v2

@@ IV Method
Given a source and target audio pair $\{x^{(0)},x^{(1)}\}$, sound morphing aims to generate intermediate sounds $x^{(\alpha(t))}$ between the audio pair given morph factors $\alpha\in[0,1]$. To account for the variation of $\alpha$ over time in Equation 1, we discretize the function $\alpha(t)$ where $t\in[0,T]$ into $N$ elements, resulting in a morphed sequence of sounds $\{x^{(\alpha_{i})}\}_{i=1}^{N}$ based on $\{\alpha_{i}\}^{N}_{i=1}$. According to the smoothness criteria proposed by [16], a desired sound morphing results should exhibit a linear perceptual change as the morph factor $\alpha$ transitions through the sequence ${\alpha_{i}}^{N}_{i=1}$. Therefore, we define $p_{i}$ to represent the perception of the morphed audio $x^{(\alpha_{i})}$ given morph factor $\alpha_{i}$. However, the relationship $\mathcal{P(\cdot)}$ between morph factor $\alpha$ and perceptual stimuli $p$ is intractable. Our objective is to determine a discrete sequence of morph factors $\{\alpha_{i}\}^{N}_{i=1}$ such that the perceptual stimulus difference $\Delta p$ remains constant for each transition. We formulate the problem as
$p_{i+1}-p_{i}\equiv\mathcal{P}(x^{(\alpha_{i+1})})-\mathcal{P}(x^{(\alpha_{i})})=\Delta p~{}$ (4)
where $i\in[1,...,N-1]$.
@@
This formulation is a refined sound morphing problem where, rather than controlling morph factor $\alpha$, we control the constant perception difference $\Delta p$ to find the optimal trajectory with morph factors $\{\alpha_{i}\}^{N}_{i=1}$ that will achieve perceptually uniform sound morphing. The overall pipeline of SoundMorpher is presented in Fig. 1 and Algorithm 1.
Algorithm 1 Overall pipeline of SoundMorpher
0:  A pre-trained AudioLDM2 pipeline including a pre-trained VAE with a encoder $g_{\theta}$ and a decoder $g_{\phi}$, a pre-trained latent diffusion model $\epsilon_{\theta}$, a pre-trained T5 model $f_{\phi}$, and a pre-trained GPT-2 model $f_{\varphi}$. Learning rates $\eta_{1},\eta_{2}$. Source and target audios $x^{(0)}$ and $x^{(1)}$. An initial text prompts $y$. perceptually uniform interpolation number $N$. tolerance error for binary search $\epsilon_{tol}$. Number of training steps for text inversion for conditional embedding optimization $T_{inv}$. Number of training steps for model adaptation $T_{adapt}$. LoRA rank $r$, step number of DDIM $T$.
0:  Start morph factor $\alpha_{start}=0$, end morph factor $\alpha_{end}=1$. Start perceptual point $p_{start}=[0,1]$, and end perceptual point $p_{end}=[1,0]$.
  # Step 1: Text-guided conditional embedding optimization
  for $i$ from $1$ to $T_{inv}$ do
     Randomly sample time step $t$ and random noise $\epsilon_{t}\sim N(0,I)$.
     $E^{(0)}\leftarrow E^{(0)}-\eta_{1}\nabla_{E^{(0)}}\mathcal{L}_{\text{DPM}}(z_{0}^{(0)},E^{(0)};\theta)$.
     $E^{(1)}\leftarrow E^{(1)}-\eta_{1}\nabla_{E^{(1)}}\mathcal{L}_{\text{DPM}}(z_{0}^{(1)},E^{(1)};\theta)$.
  end for
  # Step 2: Model adaptation with LoRA.
  for $i$ from $1$ to $T_{adapt}$ do
  end for
  # Step 3: Perceptual-uniform binary search with constant SPDP increment
  Obtaining initial latent states $z_{T}^{(0)}$ and $z_{T}^{(1)}$ by Equation 3
  $p_{list}\leftarrow$ ConstantSPDP$(N,p_{start},p_{end})$
  $\alpha_{list}\leftarrow\text{BinarySeach}(\alpha_{start},\alpha_{end},p_{list},\epsilon_{tol})$
  # Interpolate embeddings and latent state
  for $\alpha$ in $\alpha_{list}$ do
     $E^{(\alpha)}\leftarrow(1-\alpha)E^{(0)}+\alpha E^{(1)}$
     for $t$ from T todo
     end for
  end for
  $x^{(\alpha)}\leftarrow\text{vocoder}(g_{\phi}(z_{0}^{(\alpha)}))$
  return  $\{x^{(\alpha)}\}_{\alpha\in\alpha_{list}}$
@@
In Section IV-A and Section IV-B, we introduce feature interpolation and model adaptation with a pre-trained AudioLDM2. This method allows high-quality intermediate morph results to be obtained by interpolating semantic conditional embeddings of source and target sounds according to morph factor $\alpha$. In Section IV-C, we explore an explicit connection $\mathcal{P(\cdot)}$ between perceptual stimuli $p$ and morph factor $\alpha$ to achieve perceptually uniform sound morphing, as in Equation 4. This approach enables us to refine the morph factors using binary search to achieve the desired perceptual differences $\Delta p$ for each transition. In Section IV-D, we provide extensions of our method on the different morphing methods discussed in Section III-A to show the advantages of perceptually uniform sound morphing.
IV-A Feature Interpolation
Similar to other sound morphing methods, SoundMorpher achieves sound morphing by interpolating features between source and target audios. In this section, we introduce a method to bridge the source and target audios based on semantic conditional embeddings in a pre-trained AudioLDM2, leading to a controllable sound morphing method by a morph factor $\alpha$.
@@ IV-A Feature Interpolation
Interpolating optimized conditional embeddings. We first introduce a text-guided conditional embedding optimization strategy under a pre-trained AudioLDM2 [42], which retrieves corresponding conditional embeddings $E$ of the given audio data. This optimization strategy enables our sound morphing method to handle diverse types of sounds without requiring the model to be retrained on specific datasets. As mentioned in Section III-B, AudioLDM2 accepts two conditional inputs: LOA feature $Y$ and text embedding $E_{T5}$. We denote $E=\{Y,E_{T5}\}$ as the overall conditional embedding inputs for AudioLDM2. The LOA feature $Y$ is an abstraction of audio data which is semantically structured, and $E_{T5}$ captures sentence-level of representations.
$\begin{split}E^{(0)}=\min_{E}\mathcal{L}_{\text{DPM}}(z_{0}^{(0)},E;\theta)\text{ and }\\ E^{(1)}=\min_{E}\mathcal{L}_{\text{DPM}}(z_{0}^{(1)},E;\theta)\end{split}$ (5)
@@
Interpolating latent state. The conditional embedding represents the conceptual abstract of audio. We also wish to smoothly morph the content of the audio pair. Following [50], we smoothly interpolate between $z_{0}^{(0)}$ and $z_{0}^{(1)}$ by spherical linear interpolation (slerp) to their starting noise $z_{T}^{(0)}$ and $z_{T}^{(1)}$ and further obtained the interpolated latent state $z_{T}^{(\alpha)}:=\frac{\text{sin}(1-\alpha)\omega}{\text{sin}\omega}z_{T}^{(0)}+\frac{\text{sin}\alpha\omega}{\text{sin}\omega}z_{T}^{(1)}$, where $\omega=\text{arccos}(\frac{z_{T}^{(0)\intercal}z_{T}^{(1)}}{||z_{T}^{(0)}||||z_{T}^{(1)}||})$. The denoised latent variable $z_{0}^{(\alpha)}$ is obtained by applying a diffusion denoising process on the interpolated starting noise $z_{T}^{(\alpha)}$ and conditioning on the interpolated conditional embedding $E^{(\alpha)}$. The final morphed audio result $x^{(\alpha)}$ is obtained from $z_{0}^{(\alpha)}$ by the VAE decoder and a vocoder.
IV-B Model adaptation
@@
Model adaptation helps to limit the degree of morphed variation by suppressing high-density regions that not related to the given inputs [52]. To enhance the output quality of source and target audios. We follow [52] and use LoRA [53] to inject a small amount of trainable parameters for efficient model adaptation. We fine-tune AudioLDM2 with LoRA trainable parameters using $z^{(0)}$ and $z^{(1)}$. Compared to vanilla fine-tuning approaches, LoRA has advantages in training efficiency with injecting small amount of trainable parameters. The model adaptation can be defined as
$\begin{split}\min_{\Delta\theta^{\prime}}\mathcal{L}_{\text{DPM}}(z_{0}^{(0)},E^{(0)};\theta+\Delta\theta^{\prime})&+\\ \min_{\Delta\theta^{\prime}}\mathcal{L}_{\text{DPM}}(z_{0}^{(1)},E^{(1)};\theta+\Delta\theta^{\prime})&\text{ s.t. rank}(\Delta\theta^{\prime})=r\end{split}$ (6)
where $r$ represents LoRA rank.
To achieve high text alignment on inference, we use another LoRA parameters $\Delta\theta_{0}$ with rank $r_{0}$ to perform bias correction
$\begin{split}\min_{\Delta\theta_{0}}\mathcal{L}_{\text{DPM}}(z_{0}^{(0)},\varnothing;\theta+\Delta\theta_{0})&+\\ \min_{\Delta\theta_{0}}\mathcal{L}_{\text{DPM}}(z_{0}^{(1)},\varnothing;\theta+\Delta\theta_{0})&\text{ s.t. rank}(\Delta\theta_{0})=r_{0}\end{split}$ (7)
Although [52] provide a heuristic suggestion for setting the LoRA rank value in the image morphing task; however, we further investigate the relationship between LoRA rank $r$ and method performance in our emprical experiment.
IV-C Perceptually Uniform Sound Morphing
Although interpolating semantic features between the source and target audios enables a smooth transition by leveraging the inherent structure of the semantic space, our ultimate goal is to explore the complex relationship between sound perception and the morph factor to further enhance the smoothness of the morphing sequence. In this section, we propose a method to further refine the morphed results by fixing the perception difference for each transition as constant as in Equation 4. Sound perceptual distance proportion (SPDP). The relationship between morph factor $\alpha$ and perceptual stimuli $p$ is intractable. Our goal is to establish an objective quantitative metric that links $p_{i}$ and $x^{(\alpha_{i})}$ as in Equation 4. This metric should satisfy two key conditions: (1) the output $p$ should increase monotonically as $\alpha$ increases; (2) it should accurately represent sound perception and audio descriptive difference between $x^{(\alpha)}$ and $\{x^{(0)},x^{(1)}\}$, ensuring a smooth transition through intermediate states.
@@ IV-C Perceptually Uniform Sound Morphing
To this end, we propose the sound perceptual distance proportion between $x^{(\alpha)}$ and $\{x^{(0)},x^{(1)}\}$. We define $p_{i}\in\mathbb{R}^{2}$ as a 2D vector to represent the perceptual proximity of $x^{(\alpha_{i})}$ to both $x^{(0)}$ and $x^{(1)}$. Instead of extracting numerous audio features through traditional signal processing techniques, we use log magnitude Mel-scaled spectrogram to capture perceptual information and descriptive aspects of audio, aligning well with the three objective criteria outlined in Section III-A. The log Mel-spectrogram [54] provides a pseudo-3D representation of audio signals, with one axis representing time and the other representing frequency on the Mel scale [55], while the values denote the magnitude of each frequency at specific time points. This visual representation enables humans to intuitively interpret audio characteristics and descriptive information through its structured depiction of temporal and spectral features. Another advantage of using Mel-spectrogram lies in the Mel filter banks, which map frequencies to equal pitch distances that correspond to how humans perceive sound [56, 57]. A logarithmic scaling on the frequency further addresses the imbalance in human perception, which is inherently logarithmic and tends to disproportionately affect low and high-frequency regions. This approach aligns with the principles of psychoacoustics, which suggest that human auditory perception is better represented on a logarithmic frequency scale [58, 59].
Denoting $x_{mel}^{(\alpha_{i})}$ as the log magnitude Mel-spectrogram of audio $x^{(\alpha_{i})}$, the SPDP point $p_{i}$ of $x^{(\alpha_{i})}$ between $x^{(0)}$ and $x^{(1)}$ given $\alpha_{i}$ is defined as
$p_{i}=$ $[\frac{||x_{mel}^{(\alpha_{i})}-x_{mel}^{(0)}||_{2}}{||x_{mel}^{(\alpha_{i})}-x_{mel}^{(0)}||_{2}+||x_{mel}^{(\alpha_{i})}-x_{mel}^{(1)}||_{2}},$ (9)
$\frac{||x_{mel}^{(\alpha_{i})}-x_{mel}^{(1)}||_{2}}{||x_{mel}^{(\alpha_{i})}-x_{mel}^{(0)}||_{2}+||x_{mel}^{(\alpha_{i})}-x_{mel}^{(1)}||_{2}}]$
Equation 9 calculates the distance proportion of how the morphed sample $x^{(\alpha_{i})}$ closes between source $x^{(0)}$ and target $x^{(1)}$, meanwhile the L2 norm measures the distance between two log Mel spectrograms. Therefore, the summation of the two elements in $p$ are always equals to 1. Furthermore, the first element of $p_{i}$ is strictly monotonically increasing with $\alpha$ while the second element is strictly monotonically decreasing, providing a clear and interpretable relationship between the parameter $\alpha$ and the perceptual proximity of $x^{(\alpha_{i})}$ to $x^{(0)}$ and $x^{(1)}$. In summary, the log Mel-spectrogram effectively captures both perceptual and descriptive audio information, offering an elegant solution to satisfy the criterions of smoothness, correspondence, and intermediateness proposed by [16].
@@
Binary search with constant SPDP increment. To produce a perceptually smooth morphing trajectory with a constant perceptual stimuli increment, we use binary search to search for the corresponding $\{\alpha_{i}\}^{N}_{i=1}$ based on a constant $\Delta p$ as in Algorithm 2. The target SPDP sequence $\{p_{i}\}_{i=1}^{N}$ is obtained by interpolation $p_{i}=(1-\frac{i-1}{N-1})p^{(0)}+\frac{i-1}{N-1}p^{(1)}$, where the two endpoints are $p^{(0)}=[0,1]^{T}$ and $p^{(1)}=[1,0]^{T}$.
Algorithm 2 Binary search with constant $\Delta p$
0:  $\alpha_{start}$: start alpha value; $\alpha_{end}$: end alpha value; $N$: interpolate number; source audio $x^{(0)}$; target audio $x^{(1)}$;
0:  $p_{list}=[]$, $p_{start}=[0,1]^{T}$, $p_{end}=[1,0]^{T}$; $\alpha_{list}=[]$, $\alpha_{start}=0$, $\alpha_{end}=1$;
  # Step 1: Obtain target SPDP points with constant $\Delta p$.
  Procedure ConstantSPDP$(N,p_{start},p_{end})$
  for $i$ from $1$ to $N-1$ do
     $t\leftarrow\frac{i}{N-1}$
     $p_{i}\leftarrow(1-t)\times p_{start}+t\times p_{end}$
     $p_{list}\leftarrow p_{list}\cup[p_{i}]$
  end for
  # Step 2: Perform binary search given target SPDP points.
  Procedure BinarySearch($\alpha_{start},\alpha_{end},x^{(0)},x^{(1)},p_{list},\epsilon_{tol}$)
  $\alpha_{list}\leftarrow[\alpha_{start}]$, $\alpha_{cur}\leftarrow\alpha_{start}$;
  for $p_{i}$ from $p_{1}$ to $p_{N-2}$ do
     $p_{target}\leftarrow p_{i}$
     $\alpha_{t1}\leftarrow\alpha_{cur}$, $\alpha_{t2}\leftarrow\alpha_{end}$
     $\alpha_{mid}\leftarrow\frac{\alpha_{t1}+\alpha_{t2}}{2}$
     $p_{mid}\leftarrow SPDP(x^{\alpha_{mid}},x^{(0)},x^{(1)})$
     while $|p_{mid}-p_{i}|>\epsilon_{tol}$ do
        if $p_{mid}>p_{target}$ then
           $\alpha_{t2}\leftarrow\frac{\alpha_{t1}+\alpha_{t2}}{2}$
        else
           $\alpha_{t1}\leftarrow\frac{\alpha_{t1}+\alpha_{t2}}{2}$
        end if
        $\alpha_{mid}\leftarrow\frac{\alpha_{t1}+\alpha_{t2}}{2}$
        $p_{mid}\leftarrow SPDP(x^{\alpha_{mid}},x^{(0)},x^{(1)})$
     end while
     $\alpha_{list}\leftarrow\alpha_{list}\cup[\alpha_{mid}]$
  end for
  return  $\alpha_{list}$
IV-D Controllable Sound Morphing with Discrete $\alpha$ Series
Once a perceptually uniform discrete morph factor sequence $\{\alpha_{i}\}^{N}_{i=1}$ is obtained, we can further extend it to different sound morphing methods discussed in Section III-A.